OpenStudy (anonymous):

I need help in finding the enthalpy of different blends of octane and ethanol. I know how to find it for 100% ethanol and octane but i am not sure about how to find the blends. By blends i mean 0.1 mL of ethanol: 0.9 mL of octane...... 0.8 mL of ethanol, 0.2mL of octane... etc (PLEASE HELPPPP!!!)

OpenStudy (anonymous):

@chmvijay

OpenStudy (chmvijay):

can u tell me for one mole of octane and 1 mole of ethanol the enthalapy of energy ?

OpenStudy (anonymous):

for 1 mole of octane it was -5048 kJ/mol... and for 1mole of ethanol it was -1259 k/mol

OpenStudy (chmvijay):

0.1 mL of ethanol: 0.9 mL of octane is it mean 0.1moleof ethanol and 0.9 moles of octane ? i confused with ur mL?

OpenStudy (anonymous):

yes, the moles will work just the same

OpenStudy (chmvijay):

= moles of ethanol *(-1259) + moles of octane *(-5048) =0.1*(-1259) + 0.9 *(-5048) =-4532.2-125.9 = 4669 KJ mole similarly u do for other blending stuffs the final value shows that blending reduces the energy released per mole of liquid :)

OpenStudy (chmvijay):

u got it @adriana_98

OpenStudy (anonymous):

Yes, i do. Thanks very much for all your help. I really appreciate it :)

OpenStudy (chmvijay):

you are welcome :)

OpenStudy (anonymous):

Sorry for all of the question

OpenStudy (chmvijay):

hey this blog is for asking question and to get helps :) dont say sorry for this all ;) it s my pleasure that i helped you out and u learnt something from my knowledge :)

OpenStudy (anonymous):

Thanks very much. If i have any other questions,i will be sure to ask you :)..... well, bye

OpenStudy (chmvijay):

bbye :)